OpenStudy (anonymous):

This exercise requires you to enter a mathematical expression. Solve for P. pv/v=nRT/v

Nnesha (nnesha):

\[\huge\rm \frac{ pv }{ v }=\frac{ NRT }{ v }\] and you don't need a calculator

Nnesha (nnesha):

pv over v = ?

OpenStudy (anonymous):

the thing is i dont get it

Nnesha (nnesha):

you have to solve that equation for *P* so our job is to cancel out v from left side to do that we need to divide both sides by v

Nnesha (nnesha):

pv/v = ? 2/2= 1 so pv/v = ?

Nnesha (nnesha):

p is multiplying with v that's we are dividing (opposite of multiplication ) to cancel out v from left side

OpenStudy (anonymous):

does the p by itself?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Do you still not understand?

OpenStudy (anonymous):

no

Nnesha (nnesha):

yes so when you divide pv/v \[\frac{ p\cancel{v} }{\cancel{ v} }= p\]

OpenStudy (anonymous):

k then what

OpenStudy (anonymous):

Ok so you understand why she had pv/v=nrt/v

OpenStudy (anonymous):

i knew that i p is by it self .and i dont know want to do next

Nnesha (nnesha):

then nothing you need p itself and you got that.

Nnesha (nnesha):

that's it!

Nnesha (nnesha):

and i also mentioned why we dividing by v

OpenStudy (anonymous):

i dont put nRt/v

Nnesha (nnesha):

what do you mean ?

OpenStudy (anonymous):

that is put of the question

OpenStudy (anonymous):

It's not going to be a number answer

OpenStudy (anonymous):

i know

Nnesha (nnesha):

so P = nrt/v

Nnesha (nnesha):

isn't it right ?

OpenStudy (anonymous):

That's what I got

OpenStudy (anonymous):

Enter '1.230e3' in the answer box below. (Round your answer to four significant figures)

Nnesha (nnesha):

enter where ? and that's not up there ^^ your question is **This exercise requires you to enter a mathematical expression. Solve for P. pv/v=nRT/v ***

OpenStudy (anonymous):

so what is the right answer

OpenStudy (anonymous):

hello

Nnesha (nnesha):

what did you get after dividing both sides by v ?

OpenStudy (anonymous):

nothing it told it was wrong

Nnesha (nnesha):

subject name ? what's the statement ?

Nnesha (nnesha):

i guess they are asking totally different thing .....

OpenStudy (anonymous):

You can also enter your answer directly instead of using the palette. Try typing 'nRT/V' in the answer box. Notice how the expression is displayed in the answer box as you enter it. Some questions will display additional instruction on how your answer should be entered. Here instructions might be added if you need to substitute numeric values for R or which variables should only be entered in the answer box.

OpenStudy (anonymous):

it goes with pv/v=nRT/v

Nnesha (nnesha):

It's \[\huge\rm 'Pv=nRT'\] should mentioned that *'* anyway i was doing algebra stuff

Nnesha (nnesha):

and i'm sorry i don't know how to solve that for P

OpenStudy (joannablackwelder):

Were you given values for n, T, and V?